OpenStudy (anonymous):

Using the Pythagorean theorem, find the measure of the missing leg if a = 8m and c = 17m on a right triangle,

OpenStudy (maheshmeghwal9):

Pythagoras theorem states that\[\color{blue}{h^2=p^2+b^2 \implies h=\sqrt{p^2+b^2}.}\] 

OpenStudy (maheshmeghwal9):